A sphere of radius r moving with a velocity v under streamline conditions in a viscous fluid experiences a retarding force, F, given by F=krv, Where K is a constant. Derive the S.I unit and the physical dimension of k
1
Expert's answer
2021-09-30T10:20:01-0400

We use the formula and we start to substitute all the units for F (Newton), r (meters) and velocity v (meters/second) to find:


"F=krv 
\\ \implies [N]=k[m][m/s ]=k[m^2/s]
\\ \implies k=\cfrac{[N]}{[m^2/s]}=\cfrac{[Ns]}{[m^2]}=\cfrac{[(kgm/s^2)s]}{[m^2]}
\\ \implies k=\bigg[\cfrac{kg}{ms} \bigg]"


In conclusion, the SI units for the constant k are "kg/(m\cdot s)" or "Pa\cdot s".

The dimensions of k will be "\dfrac{[Mass]}{[Length*Time]}".



Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.